# Towards Massive Interaction with Generalist Robotics: A Systematic Review of XR-enabled Remote Human-Robot Interaction Systems

### v2

@@ Appendix A Data Extraction List for Included Publication
Grabbing/Picking/Placement
(Xu et al., 2022b)(Ponomareva et al., 2021)(Wei et al., 2021)(Vu et al., 2022)(Cruz Ulloa et al., 2022)(Chen et al., 2017)(Peppoloni et al., 2015b)(Kuo et al., 2021)(Yun and Jun, 2022)(Bai et al., 2018)(Omarali et al., 2020)(Xu et al., 2022a)(Jang et al., 2021b)(Hernández et al., 2020)(Moniri et al., 2016)(Hönig et al., 2015)(Brizzi et al., 2018)(Bian et al., 2018)(McHenry et al., 2021)(Xue et al., 2020)(Zaman et al., 2023)(Fan et al., 2023)(Steinke et al., 2023)(Kanazawa et al., 2023)(Meng et al., 2023)(Nenna et al., 2023)(Sanfilippo et al., 2022)
Navigation
(Chen et al., 2022)(Stedman et al., 2022)(Wibowo et al., 2021)(Stotko et al., 2019)(Ai et al., 2016)(Codd-Downey et al., 2014)(Phan et al., 2018)(Jang et al., 2021a)(Mu et al., 2021)(Theofilis et al., 2016)(Betancourt et al., 2022)(Gai et al., 2020)(Chacko et al., 2020)(Walker et al., 2019)(Grzeskowiak et al., 2020)(Wang et al., 2017)(Zhang, 2018)(Xia et al., 2023)
Industrial/Manufacturing
(Mourtzis et al., 2019)(Lipton et al., 2017)(Su et al., 2021)(Vagvolgyi et al., 2018)(Zhou et al., 2020)(Mourtzis et al., 2017)(Wang et al., 2019a)(Yew et al., 2017)(Galambos et al., 2015)(Xie et al., 2022)(Fani et al., 2018b)(Xu et al., 2018)(Wang et al., 2021b)(Wang et al., 2019b)(Nagy et al., 2019)(Aschenbrenner et al., 2018)(Rivera-Pinto et al., 2023)(Du et al., 2022a)
Multiple
(Park et al., 2018)(Rastogi and Kumar Srivastava, 2019)(Cichon and Roßmann, 2018)(Peppoloni et al., 2015a)(Du et al., 2022b)(Du et al., 2019)
No
(Gammieri et al., 2017)(Wang et al., 2019c)(Cousins et al., 2017)(Xue Er Shamaine et al., 2020)(Choi et al., 2022)(Nandhini et al., 2023)(Trinitatova and Tsetserukou, 2023)
Environment scan
(Walker et al., 2021)(Chen et al., 2020)(Zhao et al., 2017)
Surgery/Healthcare
(Zinchenko and Song, 2021)(Trejo and Hu, 2018)(Black et al., 2023)
Search
(Kalinov et al., 2021)
Game/Entertainment/Social
(Schwarz et al., 2021)(Sun et al., 2021)(Fuchino et al., 2023)
Other
(Zein et al., 2021)(Kulikovskiy et al., 2021)(Vempati et al., 2019)(Takacs et al., 2015)(Le et al., 2020)(Gong et al., 2017)(Ihara et al., 2023)(Sakashita et al., 2023)(Kim et al., 2023)(Szczurek et al., 2023)(Abdulsalam and Aouf, 2023)(Liu et al., 2023)
@@
VE(Virtual environment)
(Cruz Ulloa et al., 2022)(Walker et al., 2021)(Zein et al., 2021)(Wang et al., 2019a)(Galambos et al., 2015)(Vempati et al., 2019)(Takacs et al., 2015)(Trejo and Hu, 2018)(Cousins et al., 2017)(Hönig et al., 2015)(Wang et al., 2021b)(Grzeskowiak et al., 2020)(Wang et al., 2019b)(Gong et al., 2017)(Steinke et al., 2023)(Kanazawa et al., 2023)(Meng et al., 2023)(Liu et al., 2023)(Kuo et al., 2021)(Zinchenko and Song, 2021)(Xu et al., 2022a)(Jang et al., 2021a)(Moniri et al., 2016)(Xue Er Shamaine et al., 2020)(Brizzi et al., 2018)(Choi et al., 2022)(Sakashita et al., 2023)(Zaman et al., 2023)(Szczurek et al., 2023)(Fan et al., 2023)(Fuchino et al., 2023)
@@
User
(Ponomareva et al., 2021)(Chen et al., 2017)(Park et al., 2018)(Rastogi and Kumar Srivastava, 2019)(Schwarz et al., 2021)(Fani et al., 2018b)(Sun et al., 2021)(Du et al., 2022b)(Kim et al., 2023)(Xia et al., 2023)(Su et al., 2021)(Yun and Jun, 2022)(Moniri et al., 2016)(Choi et al., 2022)(Fan et al., 2023)(Fuchino et al., 2023)(Trinitatova and Tsetserukou, 2023)(Rivera-Pinto et al., 2023)(Black et al., 2023)
Robot(Virtual)
(Betancourt et al., 2022)(Nagy et al., 2019)(McHenry et al., 2021)(Kuo et al., 2021)(Xu et al., 2022a)(Jang et al., 2021a)(Brizzi et al., 2018)(Szczurek et al., 2023)
Robot(Real)
(Mourtzis et al., 2019)(Mourtzis et al., 2017)(Hedayati et al., 2018)(Choi et al., 2022)(Sakashita et al., 2023)
RE(Real environment)
(Hernández et al., 2020)(Chacko et al., 2020)(Walker et al., 2019)(Ihara et al., 2023)(Hedayati et al., 2018)(Zaman et al., 2023)
Object(Virtual)
(Vu et al., 2022)(Vagvolgyi et al., 2018)(Gammieri et al., 2017)(Xue et al., 2020)(Su et al., 2021)(Yun and Jun, 2022)(Zinchenko and Song, 2021)(Xue Er Shamaine et al., 2020)(Brizzi et al., 2018)
No
(Chen et al., 2022)(Xu et al., 2022b)(Wei et al., 2021)(Stedman et al., 2022)(Lipton et al., 2017)(Wibowo et al., 2021)(Kulikovskiy et al., 2021)(Zhou et al., 2020)(Stotko et al., 2019)(Peppoloni et al., 2015b)(Chen et al., 2020)(Yew et al., 2017)(Xie et al., 2022)(Ai et al., 2016)(Cichon and Roßmann, 2018)(Codd-Downey et al., 2014)(Bai et al., 2018)(Peppoloni et al., 2015a)(Phan et al., 2018)(Omarali et al., 2020)(Jang et al., 2021b)(Wang et al., 2019c)(Mu et al., 2021)(Kalinov et al., 2021)(Theofilis et al., 2016)(Zhao et al., 2017)(Gai et al., 2020)(Xu et al., 2018)(Le et al., 2020)(Hormaza et al., 2019)(Wang et al., 2017)
@@
Voice
(Du et al., 2022b)(Rivera-Pinto et al., 2023)
Video
(Steinke et al., 2023)(Meng et al., 2023)(Hedayati et al., 2018)(Zinchenko and Song, 2021)(Xu et al., 2022a)(Zaman et al., 2023)(Szczurek et al., 2023)(Fan et al., 2023)(Kanazawa et al., 2023)
Text
(Wang et al., 2019a)(Wang et al., 2021b)(Wang et al., 2019b)(Mourtzis et al., 2019)(Kuo et al., 2021)(Yun and Jun, 2022)(Xu et al., 2022a)(Chacko et al., 2020)(Xue Er Shamaine et al., 2020)(Szczurek et al., 2023)(Kanazawa et al., 2023)
Ray
(Kuo et al., 2021)(Moniri et al., 2016)(Brizzi et al., 2018)(Choi et al., 2022)(Zaman et al., 2023)
Highlight
(Betancourt et al., 2022)(Xue et al., 2020)(Xu et al., 2022a)(Jang et al., 2021a)(Xue Er Shamaine et al., 2020)(Brizzi et al., 2018)(Szczurek et al., 2023)
Haptic
(Ponomareva et al., 2021)(Chen et al., 2017)(Park et al., 2018)(Rastogi and Kumar Srivastava, 2019)(Schwarz et al., 2021)(Fani et al., 2018b)(Sun et al., 2021)(Kim et al., 2023)(Xia et al., 2023)(Su et al., 2021)(Yun and Jun, 2022)(Du et al., 2022b)(Fan et al., 2023)(Fuchino et al., 2023)(Trinitatova and Tsetserukou, 2023)(Black et al., 2023)
Graphic
(Vagvolgyi et al., 2018)(Gammieri et al., 2017)(Vempati et al., 2019)(Trejo and Hu, 2018)(Gong et al., 2017)(Vu et al., 2022)(Zein et al., 2021)(Zinchenko and Song, 2021)(Xue Er Shamaine et al., 2020)(Brizzi et al., 2018)(Ihara et al., 2023)(Sakashita et al., 2023)(Zaman et al., 2023)(Trinitatova and Tsetserukou, 2023)
Avatar
(Galambos et al., 2015)(Takacs et al., 2015)(Cousins et al., 2017)(Hönig et al., 2015)(Grzeskowiak et al., 2020)(Liu et al., 2023)(Jang et al., 2021a)(Moniri et al., 2016)(Choi et al., 2022)(Ihara et al., 2023)(Sakashita et al., 2023)(Zaman et al., 2023)
3D Object
(Cruz Ulloa et al., 2022)(Walker et al., 2021)(Mourtzis et al., 2017)(Hernández et al., 2020)(Walker et al., 2019)(Nagy et al., 2019)(McHenry et al., 2021)(Vu et al., 2022)(Mourtzis et al., 2019)(Su et al., 2021)(Zein et al., 2021)(Hedayati et al., 2018)(Jang et al., 2021a)(Chacko et al., 2020)(Choi et al., 2022)(Fuchino et al., 2023)(Trinitatova and Tsetserukou, 2023)(Rivera-Pinto et al., 2023)(Black et al., 2023)
No
(Chen et al., 2022)(Xu et al., 2022b)(Wei et al., 2021)(Stedman et al., 2022)(Lipton et al., 2017)(Wibowo et al., 2021)(Kulikovskiy et al., 2021)(Zhou et al., 2020)(Stotko et al., 2019)(Peppoloni et al., 2015b)(Chen et al., 2020)(Yew et al., 2017)(Xie et al., 2022)(Ai et al., 2016)(Cichon and Roßmann, 2018)(Codd-Downey et al., 2014)(Bai et al., 2018)(Peppoloni et al., 2015a)(Phan et al., 2018)(Omarali et al., 2020)(Jang et al., 2021b)(Wang et al., 2019c)(Mu et al., 2021)(Kalinov et al., 2021)(Theofilis et al., 2016)(Zhao et al., 2017)(Gai et al., 2020)(Xu et al., 2018)(Le et al., 2020)(Hormaza et al., 2019)(Wang et al., 2017)
@@
N/A
(Stedman et al., 2022)(Walker et al., 2021)(Zhou et al., 2020)(Mourtzis et al., 2017)(Rastogi and Kumar Srivastava, 2019)(Yew et al., 2017)(Galambos et al., 2015)(Gammieri et al., 2017)(Xie et al., 2022)(Cichon and Roßmann, 2018)(Zinchenko and Song, 2021)(Codd-Downey et al., 2014)(Bai et al., 2018)(Takacs et al., 2015)(Phan et al., 2018)(Xu et al., 2022a)(Jang et al., 2021b)(Wang et al., 2019c)(Mu et al., 2021)(Theofilis et al., 2016)(Hernández et al., 2020)(Zhao et al., 2017)(Moniri et al., 2016)(Xu et al., 2018)(Sun et al., 2021)(Hönig et al., 2015)(Hormaza et al., 2019)(Xue Er Shamaine et al., 2020)(Wang et al., 2017)(Zhang, 2018)
@@
Time/accuracy of the task
(Xu et al., 2022b)(Ponomareva et al., 2021)(Cruz Ulloa et al., 2022)(Mourtzis et al., 2019)(Zein et al., 2021)(Hedayati et al., 2018)(Stotko et al., 2019)(Fani et al., 2018b)(Omarali et al., 2020)(Trejo and Hu, 2018)(Le et al., 2020)(Walker et al., 2019)(Zaman et al., 2023)(Kim et al., 2023)(Kanazawa et al., 2023)(Trinitatova and Tsetserukou, 2023)(Meng et al., 2023)(Nenna et al., 2023)(Xia et al., 2023)(Liu et al., 2023)(Black et al., 2023)(Wei et al., 2021)(Vu et al., 2022)(Lipton et al., 2017)(Su et al., 2021)(Wibowo et al., 2021)(Vagvolgyi et al., 2018)(Kulikovskiy et al., 2021)(Chen et al., 2017)(Park et al., 2018)(Peppoloni et al., 2015b)
@@
Questionnaire
(Ponomareva et al., 2021)(Cruz Ulloa et al., 2022)(Mourtzis et al., 2019)(Zein et al., 2021)(Hedayati et al., 2018)(Stotko et al., 2019)(Schwarz et al., 2021)(Fani et al., 2018b)(Omarali et al., 2020)(Kalinov et al., 2021)(Trejo and Hu, 2018)(Le et al., 2020)(Walker et al., 2019)(Ihara et al., 2023)(Zaman et al., 2023)(Kim et al., 2023)(Kanazawa et al., 2023)(Trinitatova and Tsetserukou, 2023)(Meng et al., 2023)(Nenna et al., 2023)(Xia et al., 2023)(Liu et al., 2023)(Rivera-Pinto et al., 2023)
Comparison
(Xu et al., 2022b)(Yun and Jun, 2022)(Betancourt et al., 2022)(Cousins et al., 2017)(Abdulsalam and Aouf, 2023)(Trinitatova and Tsetserukou, 2023)(Xia et al., 2023)
Interview
(Hedayati et al., 2018)(Walker et al., 2019)(Ihara et al., 2023)(Sanfilippo et al., 2022)
AR/VR
(Chen et al., 2020)(Black et al., 2023)
Other
(Chen et al., 2020)(Yun and Jun, 2022)(Ai et al., 2016)(Gai et al., 2020)
@@
Multi-user - One robot
(Galambos et al., 2015)
One user - Multi-robot
(Jang et al., 2021a)(Gong et al., 2017)
Multi-user - Multi-robot
(Phan et al., 2018)(Hönig et al., 2015)(Aschenbrenner et al., 2018)
One-user - One-user+One robot
(Mourtzis et al., 2019)(Mourtzis et al., 2017)(Schwarz et al., 2021)(Moniri et al., 2016)(Sakashita et al., 2023)(Zaman et al., 2023)(Szczurek et al., 2023)(Fuchino et al., 2023)(Liu et al., 2023)(Black et al., 2023)
One-user - Multi-user+One robot
(Walker et al., 2021)

### v3

@@ Appendix A Data Extraction List for Included Publication
Grabbing/Picking/Placement (Xu et al., 2022b)(Ponomareva et al., 2021)(Wei et al., 2021)(Vu et al., 2022)(Cruz Ulloa et al., 2022)(Chen et al., 2017)(Peppoloni et al., 2015b)(Kuo et al., 2021)(Yun and Jun, 2022)(Bai et al., 2018)(Omarali et al., 2020)(Xu et al., 2022a)(Jang et al., 2021b)(Hernández et al., 2020)(Moniri et al., 2016)(Hönig et al., 2015)(Brizzi et al., 2018)(Bian et al., 2018)(McHenry et al., 2021)(Xue et al., 2020)(Zaman et al., 2023)(Fan et al., 2023)(Steinke et al., 2023)(Kanazawa et al., 2023)(Meng et al., 2023)(Nenna et al., 2023)(Sanfilippo et al., 2022)
Navigation (Chen et al., 2022)(Stedman et al., 2022)(Wibowo et al., 2021)(Stotko et al., 2019)(Ai et al., 2016)(Codd-Downey et al., 2014)(Phan et al., 2018)(Jang et al., 2021a)(Mu et al., 2021)(Theofilis et al., 2016)(Betancourt et al., 2022)(Gai et al., 2020)(Chacko et al., 2020)(Walker et al., 2019)(Grzeskowiak et al., 2020)(Wang et al., 2017)(Zhang, 2018)(Xia et al., 2023)
Industrial/Manufacturing (Mourtzis et al., 2019)(Lipton et al., 2017)(Su et al., 2021)(Vagvolgyi et al., 2018)(Zhou et al., 2020)(Mourtzis et al., 2017)(Wang et al., 2019a)(Yew et al., 2017)(Galambos et al., 2015)(Xie et al., 2022)(Fani et al., 2018b)(Xu et al., 2018)(Wang et al., 2021b)(Wang et al., 2019b)(Nagy et al., 2019)(Aschenbrenner et al., 2018)(Rivera-Pinto et al., 2023)(Du et al., 2022a)
Multiple (Park et al., 2018)(Rastogi and Kumar Srivastava, 2019)(Cichon and Roßmann, 2018)(Peppoloni et al., 2015a)(Du et al., 2022b)(Du et al., 2019)
No (Gammieri et al., 2017)(Wang et al., 2019c)(Cousins et al., 2017)(Xue Er Shamaine et al., 2020)(Choi et al., 2022)(Nandhini et al., 2023)(Trinitatova and Tsetserukou, 2023)
Environment scan (Walker et al., 2021)(Chen et al., 2020)(Zhao et al., 2017)
Surgery/Healthcare (Zinchenko and Song, 2021)(Trejo and Hu, 2018)(Black et al., 2023)
Search (Kalinov et al., 2021)
Game/Entertainment/Social (Schwarz et al., 2021)(Sun et al., 2021)(Fuchino et al., 2023)
Other (Zein et al., 2021)(Kulikovskiy et al., 2021)(Vempati et al., 2019)(Takacs et al., 2015)(Le et al., 2020)(Gong et al., 2017)(Ihara et al., 2023)(Sakashita et al., 2023)(Kim et al., 2023)(Szczurek et al., 2023)(Abdulsalam and Aouf, 2023)(Liu et al., 2023)
@@
VE(Virtual environment) (Cruz Ulloa et al., 2022)(Walker et al., 2021)(Zein et al., 2021)(Wang et al., 2019a)(Galambos et al., 2015)(Vempati et al., 2019)(Takacs et al., 2015)(Trejo and Hu, 2018)(Cousins et al., 2017)(Hönig et al., 2015)(Wang et al., 2021b)(Grzeskowiak et al., 2020)(Wang et al., 2019b)(Gong et al., 2017)(Steinke et al., 2023)(Kanazawa et al., 2023)(Meng et al., 2023)(Liu et al., 2023)(Kuo et al., 2021)(Zinchenko and Song, 2021)(Xu et al., 2022a)(Jang et al., 2021a)(Moniri et al., 2016)(Xue Er Shamaine et al., 2020)(Brizzi et al., 2018)(Choi et al., 2022)(Sakashita et al., 2023)(Zaman et al., 2023)(Szczurek et al., 2023)(Fan et al., 2023)(Fuchino et al., 2023)
@@
User (Ponomareva et al., 2021)(Chen et al., 2017)(Park et al., 2018)(Rastogi and Kumar Srivastava, 2019)(Schwarz et al., 2021)(Fani et al., 2018b)(Sun et al., 2021)(Du et al., 2022b)(Kim et al., 2023)(Xia et al., 2023)(Su et al., 2021)(Yun and Jun, 2022)(Moniri et al., 2016)(Choi et al., 2022)(Fan et al., 2023)(Fuchino et al., 2023)(Trinitatova and Tsetserukou, 2023)(Rivera-Pinto et al., 2023)(Black et al., 2023)
Robot(Virtual) (Betancourt et al., 2022)(Nagy et al., 2019)(McHenry et al., 2021)(Kuo et al., 2021)(Xu et al., 2022a)(Jang et al., 2021a)(Brizzi et al., 2018)(Szczurek et al., 2023)
Robot(Real) (Mourtzis et al., 2019)(Mourtzis et al., 2017)(Hedayati et al., 2018)(Choi et al., 2022)(Sakashita et al., 2023)
RE(Real environment) (Hernández et al., 2020)(Chacko et al., 2020)(Walker et al., 2019)(Ihara et al., 2023)(Hedayati et al., 2018)(Zaman et al., 2023)
Object(Virtual) (Vu et al., 2022)(Vagvolgyi et al., 2018)(Gammieri et al., 2017)(Xue et al., 2020)(Su et al., 2021)(Yun and Jun, 2022)(Zinchenko and Song, 2021)(Xue Er Shamaine et al., 2020)(Brizzi et al., 2018)
No (Chen et al., 2022)(Xu et al., 2022b)(Wei et al., 2021)(Stedman et al., 2022)(Lipton et al., 2017)(Wibowo et al., 2021)(Kulikovskiy et al., 2021)(Zhou et al., 2020)(Stotko et al., 2019)(Peppoloni et al., 2015b)(Chen et al., 2020)(Yew et al., 2017)(Xie et al., 2022)(Ai et al., 2016)(Cichon and Roßmann, 2018)(Codd-Downey et al., 2014)(Bai et al., 2018)(Peppoloni et al., 2015a)(Phan et al., 2018)(Omarali et al., 2020)(Jang et al., 2021b)(Wang et al., 2019c)(Mu et al., 2021)(Kalinov et al., 2021)(Theofilis et al., 2016)(Zhao et al., 2017)(Gai et al., 2020)(Xu et al., 2018)(Le et al., 2020)(Hormaza et al., 2019)(Wang et al., 2017)
@@
Voice (Du et al., 2022b)(Rivera-Pinto et al., 2023)
Video (Steinke et al., 2023)(Meng et al., 2023)(Hedayati et al., 2018)(Zinchenko and Song, 2021)(Xu et al., 2022a)(Zaman et al., 2023)(Szczurek et al., 2023)(Fan et al., 2023)(Kanazawa et al., 2023)
Text (Wang et al., 2019a)(Wang et al., 2021b)(Wang et al., 2019b)(Mourtzis et al., 2019)(Kuo et al., 2021)(Yun and Jun, 2022)(Xu et al., 2022a)(Chacko et al., 2020)(Xue Er Shamaine et al., 2020)(Szczurek et al., 2023)(Kanazawa et al., 2023)
Ray (Kuo et al., 2021)(Moniri et al., 2016)(Brizzi et al., 2018)(Choi et al., 2022)(Zaman et al., 2023)
Highlight (Betancourt et al., 2022)(Xue et al., 2020)(Xu et al., 2022a)(Jang et al., 2021a)(Xue Er Shamaine et al., 2020)(Brizzi et al., 2018)(Szczurek et al., 2023)
Haptic (Ponomareva et al., 2021)(Chen et al., 2017)(Park et al., 2018)(Rastogi and Kumar Srivastava, 2019)(Schwarz et al., 2021)(Fani et al., 2018b)(Sun et al., 2021)(Kim et al., 2023)(Xia et al., 2023)(Su et al., 2021)(Yun and Jun, 2022)(Du et al., 2022b)(Fan et al., 2023)(Fuchino et al., 2023)(Trinitatova and Tsetserukou, 2023)(Black et al., 2023)
Graphic (Vagvolgyi et al., 2018)(Gammieri et al., 2017)(Vempati et al., 2019)(Trejo and Hu, 2018)(Gong et al., 2017)(Vu et al., 2022)(Zein et al., 2021)(Zinchenko and Song, 2021)(Xue Er Shamaine et al., 2020)(Brizzi et al., 2018)(Ihara et al., 2023)(Sakashita et al., 2023)(Zaman et al., 2023)(Trinitatova and Tsetserukou, 2023)
Avatar (Galambos et al., 2015)(Takacs et al., 2015)(Cousins et al., 2017)(Hönig et al., 2015)(Grzeskowiak et al., 2020)(Liu et al., 2023)(Jang et al., 2021a)(Moniri et al., 2016)(Choi et al., 2022)(Ihara et al., 2023)(Sakashita et al., 2023)(Zaman et al., 2023)
3D Object (Cruz Ulloa et al., 2022)(Walker et al., 2021)(Mourtzis et al., 2017)(Hernández et al., 2020)(Walker et al., 2019)(Nagy et al., 2019)(McHenry et al., 2021)(Vu et al., 2022)(Mourtzis et al., 2019)(Su et al., 2021)(Zein et al., 2021)(Hedayati et al., 2018)(Jang et al., 2021a)(Chacko et al., 2020)(Choi et al., 2022)(Fuchino et al., 2023)(Trinitatova and Tsetserukou, 2023)(Rivera-Pinto et al., 2023)(Black et al., 2023)
No (Chen et al., 2022)(Xu et al., 2022b)(Wei et al., 2021)(Stedman et al., 2022)(Lipton et al., 2017)(Wibowo et al., 2021)(Kulikovskiy et al., 2021)(Zhou et al., 2020)(Stotko et al., 2019)(Peppoloni et al., 2015b)(Chen et al., 2020)(Yew et al., 2017)(Xie et al., 2022)(Ai et al., 2016)(Cichon and Roßmann, 2018)(Codd-Downey et al., 2014)(Bai et al., 2018)(Peppoloni et al., 2015a)(Phan et al., 2018)(Omarali et al., 2020)(Jang et al., 2021b)(Wang et al., 2019c)(Mu et al., 2021)(Kalinov et al., 2021)(Theofilis et al., 2016)(Zhao et al., 2017)(Gai et al., 2020)(Xu et al., 2018)(Le et al., 2020)(Hormaza et al., 2019)(Wang et al., 2017)
@@
N/A (Stedman et al., 2022)(Walker et al., 2021)(Zhou et al., 2020)(Mourtzis et al., 2017)(Rastogi and Kumar Srivastava, 2019)(Yew et al., 2017)(Galambos et al., 2015)(Gammieri et al., 2017)(Xie et al., 2022)(Cichon and Roßmann, 2018)(Zinchenko and Song, 2021)(Codd-Downey et al., 2014)(Bai et al., 2018)(Takacs et al., 2015)(Phan et al., 2018)(Xu et al., 2022a)(Jang et al., 2021b)(Wang et al., 2019c)(Mu et al., 2021)(Theofilis et al., 2016)(Hernández et al., 2020)(Zhao et al., 2017)(Moniri et al., 2016)(Xu et al., 2018)(Sun et al., 2021)(Hönig et al., 2015)(Hormaza et al., 2019)(Xue Er Shamaine et al., 2020)(Wang et al., 2017)(Zhang, 2018)
@@
Time/accuracy of the task (Xu et al., 2022b)(Ponomareva et al., 2021)(Cruz Ulloa et al., 2022)(Mourtzis et al., 2019)(Zein et al., 2021)(Hedayati et al., 2018)(Stotko et al., 2019)(Fani et al., 2018b)(Omarali et al., 2020)(Trejo and Hu, 2018)(Le et al., 2020)(Walker et al., 2019)(Zaman et al., 2023)(Kim et al., 2023)(Kanazawa et al., 2023)(Trinitatova and Tsetserukou, 2023)(Meng et al., 2023)(Nenna et al., 2023)(Xia et al., 2023)(Liu et al., 2023)(Black et al., 2023)(Wei et al., 2021)(Vu et al., 2022)(Lipton et al., 2017)(Su et al., 2021)(Wibowo et al., 2021)(Vagvolgyi et al., 2018)(Kulikovskiy et al., 2021)(Chen et al., 2017)(Park et al., 2018)(Peppoloni et al., 2015b)
@@
Questionnaire (Ponomareva et al., 2021)(Cruz Ulloa et al., 2022)(Mourtzis et al., 2019)(Zein et al., 2021)(Hedayati et al., 2018)(Stotko et al., 2019)(Schwarz et al., 2021)(Fani et al., 2018b)(Omarali et al., 2020)(Kalinov et al., 2021)(Trejo and Hu, 2018)(Le et al., 2020)(Walker et al., 2019)(Ihara et al., 2023)(Zaman et al., 2023)(Kim et al., 2023)(Kanazawa et al., 2023)(Trinitatova and Tsetserukou, 2023)(Meng et al., 2023)(Nenna et al., 2023)(Xia et al., 2023)(Liu et al., 2023)(Rivera-Pinto et al., 2023)
Comparison (Xu et al., 2022b)(Yun and Jun, 2022)(Betancourt et al., 2022)(Cousins et al., 2017)(Abdulsalam and Aouf, 2023)(Trinitatova and Tsetserukou, 2023)(Xia et al., 2023)
Interview (Hedayati et al., 2018)(Walker et al., 2019)(Ihara et al., 2023)(Sanfilippo et al., 2022)
AR/VR (Chen et al., 2020)(Black et al., 2023)
Other (Chen et al., 2020)(Yun and Jun, 2022)(Ai et al., 2016)(Gai et al., 2020)
@@
Multi-user - One robot (Galambos et al., 2015)
One user - Multi-robot (Jang et al., 2021a)(Gong et al., 2017)
Multi-user - Multi-robot (Phan et al., 2018)(Hönig et al., 2015)(Aschenbrenner et al., 2018)
One-user - One-user+One robot (Mourtzis et al., 2019)(Mourtzis et al., 2017)(Schwarz et al., 2021)(Moniri et al., 2016)(Sakashita et al., 2023)(Zaman et al., 2023)(Szczurek et al., 2023)(Fuchino et al., 2023)(Liu et al., 2023)(Black et al., 2023)
One-user - Multi-user+One robot (Walker et al., 2021)